OpenStudy (anonymous):

Find the equation of the perpendicular line to each of the following that passes through the point indicated. 4x + 5y = - 40 passing through the point (-1, 3) Answer 5x - 4y = -17 5x + 4y = 17 5x + 4y = -17 -5x + 4y = 17

OpenStudy (anonymous):

you found it :D

OpenStudy (anonymous):

well first 4x + 5y = - 40 sfy to get it in y=mx+b form y = -(4/5)*x-8 m x b the m is your slope to get perpendicular you flip it and multiply by - so a slope perpendicular to it is 5/4

OpenStudy (anonymous):

-1,3 so now you can plug in the coordinants to y=mx+b 3=5/4*(-1)+b sfb b=17/4 so y=5/4 *x +(17/4) tto get it in the form the answer asks y-5/4*x=17/4 subtract 5/4x 4*y-5*x=17 multiply by 4

OpenStudy (anonymous):

rearrange the x and y terms your answer is -5x + 4y = 17

OpenStudy (anonymous):

:D Thank you so much! Dude you are awesome!

OpenStudy (anonymous):

so long as you understand :)